Jack and Mary are both in their eighties. Theyre thinking of selling their house for $600,000 and moving into an apartment complex for seniors. The senior complex will cost $60,000 per year, payable in full at the beginning of each year. This payment covers all of their costs (food, rental, entertainment, and medical) a. If they can earn 6% annually on the proceeds from their house and if they live for 10 more years, how much will they be able to leave to their children as an inheritance? b. What is the longest they can live from the apartment proceeds before the money runs out? C. Find the answers to parts a and b above assuming that the interest rate is 7% Make sure you build a model that looks good in addition to being correct. 796.
Answer #1

Solution :- (a)

Here the payment start at the beginning of year so

Calculate Present value of the payments of 10 years = Cashflow*{1+ P/A(6%,9)} = 60000*(1+6.8017) =

=60000*7.8017 = 468102

The amount receive by selling house = 600000

Therefore amount saves In today's value =

=60000 - 468102 = 131898$

The amount left in inheritance = 131898*(1+0.06)10

= 131898*1.79 = $236097.42

(B) as per condition

600000 = 60000*{1+P/A{6%,n}}

10 = 1 + P/A(6%,n)

P/A(6%,n) = 9

As per PVAF table n = 13.5

Therefore they can live maximum for 13.5+1 = 14.5years approx

This question is too long bro so please ask other as separate one